OpenStudy (anonymous):

What is the solution of the system? . y = x + 5 y = –5x – 1

sam (.sam.):

you can see both equation "y" is same, x + 5= –5x – 1

OpenStudy (anonymous):

no its not

sam (.sam.):

x+5=-5x-1 Move all terms not containing x to the right-hand side of the equation. 6x+5=-1 Move all terms not containing x to the right-hand side of the equation. 6x=-6 Divide each term in the equation by 6. x=-1

OpenStudy (anonymous):

the solution is what

sam (.sam.):

y = x + 5 y = -1 + 5 y=4

OpenStudy (anonymous):

no not in equation use a graph to find its intersection

OpenStudy (anonymous):

so its 4,-1

sam (.sam.):

(-1,4)

OpenStudy (anonymous):

wow i had that answer but thats not any multiple choice answer

sam (.sam.):